OpenStudy (anonymous):

PPPPLLEASE!!! URGEEEENTT!! THNX M is directly proportional to L^3 When L=2, M=160 Find the value of M when L=3

OpenStudy (anonymous):

please answer!!

jimthompson5910 (jim_thompson5910):

M is directly proportional to L^3 translates to M = k*L^3 where k is some fixed number

jimthompson5910 (jim_thompson5910):

we are given When L=2, M=160 so plug this in to get M = k*L^3 160 = k*2^3 and solve for k tell me what you get

OpenStudy (anonymous):

is the final answer 540?

jimthompson5910 (jim_thompson5910):

yes you should have gotten k = 20, so M = 20*L^3 when L = 3, we get M = 20*L^3 M = 20*3^3 M = 20*27 M = 540